# On-Policy Policy Gradient Reinforcement Learning Without On-Policy Sampling

### v2

@@ D.1 Bias and Variance of props
In Figure 7, we investigate the bias and variance of the empirical state-action visitation distribution $d^{\mathcal{D}}({\bm{s}},{\bm{a}})$ under props, ros, and on-policy sampling. Our experimental setup is similar to the setup in Section 6.1. We collect $T$ samples with a fixed policy that samples actions uniformly at random in each state. We compute the bias and variance averaged over all $({\bm{s}},{\bm{a}})\in{\mathcal{S}}\times{\mathcal{A}}$ computed as follows:
$\text{bias}=\frac{1}{|{\mathcal{S}}\times{\mathcal{A}}|}\sum_{({\bm{s}},{\bm{a}})\in{\mathcal{S}}\times{\mathcal{A}}}\left(\mathbb{E}\left[d^{{\mathcal{D}}}({\bm{s}},{\bm{a}})\right]-d_{\pi_{\bm{\theta}}}({\bm{s}},{\bm{a}})\right)$ (10)
$\text{variance}=\frac{1}{|{\mathcal{S}}\times{\mathcal{A}}|}\sum_{({\bm{s}},{\bm{a}})\in{\mathcal{S}}\times{\mathcal{A}}}\mathbb{E}\left[\left(d^{{\mathcal{D}}}({\bm{s}},{\bm{a}})-d_{\pi_{\bm{\theta}}}({\bm{s}},{\bm{a}})\right)^{2}\right]$ (11)
As shown in Figure 7, the visitation distribution under props and ros empirical have near zero bias (note that the vertical axis has scale $10^{-18}$) and have lower variance than on-policy sampling.
@@ D.2 Correcting Sampling Error for a Fixed Target Policy
Figure 10: Sampling error ablations with a fixed, expert target policy. Here, “no clipping” refers to setting $\epsilon_{\textsc{props}}{}{}=\infty$, and “no regularization” refers to setting $\lambda=0$. Solid curves denote the mean over 5 seeds, and shaded regions denote 95% bootstrap confidence intervals.
In this appendix, we expand upon results presented in Section 6.1 of the main paper and provide additional experiments investigating the degree to which props reduces sampling error with respect to a fixed target policy. We include empirical results for all six MuJoCo benchmark tasks as well as ablation studies investigating the effects of clipping and regularization.
We tune props and ros using a hyperparameter sweep. For props, we consider learning rates in $\{10^{-3},10^{-4}\}$, regularization coefficients $\lambda\in\{0.01,0.1,0.3\}$, and props target KLs in $\delta_{\textsc{props}}{}\in\{0.05,0.1\}$. We fix $\epsilon_{\textsc{props}}{}=0.3$ across all experiments. For ros, we consider learning rates in $\{10^{-3},10^{-4},10^{-5}\}$. We report results for the hyperparameters yielding the lowest sampling error.
Figure 8 and 9 show sampling error computed with a fixed expert and randomly initialized target policy, respectively. We see that props achieves lower sampling error than both ros and on-policy sampling across all tasks. ros shows little to no improvement over on-policy sampling, highlighting the difficulty of applying ros to higher dimensional tasks with continuous actions.
Figure 10 ablates the effects of props’s clipping mechanism and regularization on sampling error reduction. We ablate clipping by setting $\epsilon_{\textsc{props}}{}{}=\infty$, and we ablate regularization by setting $\lambda=0$. We use a fixed expert target policy and use the same tuning procedure described earlier in this appendix. In all tasks, props achieves higher sampling error without clipping nor regularization than it does with clipping and regularization. However, it nevertheless outperforms on-policy sampling in all tasks except Hopper where it matches the performance of on-policy sampling. Only including regularization slightly decreases sampling error, whereas clipping alone produces sampling error only slightly higher than that achieved by props with both regularization and clipping. These observations indicate that while regularization in is helpful, clipping has a stronger effect on sampling error reduction than regularization when the target policy is fixed.
@@ D.3 Correcting Sampling Error During RL Training
In this appendix, we include additional experiments investigating the degree to which props reduces sampling error during rl training, expanding upon results presented in Section 6.2 of the main paper. We include sampling error curves for all six MuJoCo benchmark tasks and additionally provide ablation studies investigating the effects of clipping and regularization on sampling error reduction and data efficiency in the rl setting.
As shown in Fig 11, props achieves lower sampling error than on-policy sampling throughout training in 5 out of 6 tasks. We observe that props increases sampling error but nevertheless improves data efficiency in HalfCheetah as shown in Figure 4(a). This result likely arises from our tuning procedure in which we selected hyperparameters yielding the largest return. Although lower sampling error intuitively correlates with increased data efficiency, it is nevertheless possible to achieve high return without reducing sampling error.
In our next set of experiments, we ablate the effects of props’s clipping mechanism and regularization on sampling error reduction and data efficiency. We ablate clipping by tuningrl agents with $\epsilon_{\textsc{props}}{}{}=\infty$, and we ablate regularization by tuning rl agents with $\lambda=0$. Figure 12 and Figure 13 show sampling error curves without clipping and without regularization, respectively. Without clipping, props achieves larger sampling than on-policy sampling in all tasks except Humanoid. Without regularization, props achieves larger sampling error in 3 out of 6 tasks. These observations indicate that while clipping and regularization both help reduce sampling during rl training, clipping has a stronger effect on sampling error reduction. As shown in Figure 14 props data efficiency generally decreases when we remove clipping or regularization.
Lastly, we consider training with larger buffer sizes $b$ in Figure 15. We find that data efficiency may decrease with a larger buffer size. Intuitively, the more historic data kept around, the more data that must be collected to impact the aggregate data distribution.
Figure 14: IQM return over 50 seeds of props with and without clipping or regularizing the props objective. Shaded regions denote 95% bootstrap confidence intervals.
Figure 15: IQM return over 50 seeds for props with different buffer sizes. We exclude $b=8$ for Humanoid-v4 due to the expense of training and tuning. Shaded regions denote 95% bootstrap confidence intervals.
D.4 Discrete-Action Tasks
We include 3 additional discrete-action domains of varying complexity. The first two are the widely used OpenAI gym domains CartPole-v1 and LunarLander-v2 [Brockman et al., 2016]. The third is a 2D navigation task, Discrete2D100-v0, in which the agent must reach a randomly sampled goal. There are 100 actions, each action corresponding to different directions in which the agent can move. From Figure 16 and 17 we observe that props with $b=2$ achieves larger returns than ppo and ppo-buffer all throughout training in all three tasks. props with $b=1$ (no historic data) achieves larger returns than ppo all throughout training in all three tasks and even outperforms ppo-buffer in CartPole-v1 and Discrete2D100-v0 even though ppo-buffer learns from twice as much data. Thus, props can improve data efficiency without historic data.
Figure 16: IQM return for discrete action tasks over 50 seeds. Shaded regions denote 95% bootstrap confidence intervals.
Figure 17: Performance profiles for discrete-action tasks over 50 seeds. Higher values correspond to more reliable convergence to high-return policies. Shaded regions denote 95% bootstrap confidence intervals.
@@ D.4 Discrete-Action Tasks
D.5 Runtime Comparisons
Figure 18 shows runtimes for props, ppo-buffer, and ppo averaged over 3 runs. We trained all agents on a MacBook Air with an M1 CPU and use the same tuned hyperparameters used throughout the paper. props takes at most twice as long as ppo-buffer; intuitively, both props and ppo-buffer learn from the same amount of data but props learns two policies.
We note that ppo-buffer is faster than ppo is HalfCheetah-v4 because, with our tuned hyperparameters, ppo-buffer performs fewer target policy updates than ppo. In particular, ppo-buffer is updating its target policy every 4096 steps, whereas ppo is updating the target policy every 1024 steps.
D.6 Hyperparameter Tuning for RL Training
For all rl experiments in Section 6.2 and Appendix D.3, we tune props, ppo-buffer, and ppo separately using a hyperparameter sweep over parameters listed in Table 2 and fix the hyperparameters in Table 3 across all experiments. Since we consider a wide range of hyperparameter values, we ran 10 independent training runs for each hyperparameter setting. We then performed 50 independent training runs for the hyperparameters settings yielding the largest returns at the end of rl training. We report results for these hyperparameters in the main paper. Figure 19 shows training curves obtained from a subset of our hyperparameter sweep.
ppo learning rate $10^{-3},10^{-4}$, linearly annealed to 00 over training
@@ D.6 Hyperparameter Tuning for RL Training
props behavior batch size $m$ $256,512,1024,2048,4096$ satisfying $m\leq n$
props KL cutoff $\delta_{\textsc{props}}{}$ $0.03,0.05,0.1$
props regularizer coefficient $\lambda$ $0.01,0.1,0.3$
@@
Buffer size $b$ 2 target batches (also 3, 4, and 8 in Figure 15)
@@
ppo and props optimizers Adam [Kingma and Ba, 2015]
@@
Table 3: Hyperparameters fixed across all experiments. We use the ppo implementation provided by CleanRL [Huang et al., 2022].
(a) props kl cutoff values $\delta_{\textsc{props}}{}$.
(b) Regularization coefficients $\lambda$.
(c) Behavior batch sizes $m$ (i.e. the number of steps between behavior policy updates).
Figure 19: A subset of results obtained from our hyperparameter sweep. Default hyperparameter values are as follows: PROPS kl cutoff $\delta_{\textsc{props}}{}=0.03$; regularization coefficient $\lambda=0.1$; behavior batch size $m=256$. Darker colors indicate larger hyperparameter values. Solid and dashed lines have the PROPS learning rate set to $1\cdot 10^{-3}$ and $1\cdot 10^{-4}$, respectively. Curves denote averages over 10 seeds, and shaded regions denote 95% confidence intervals.

### v3

@@ D.1 Bias and Variance of PROPS
In Fig. 8, we investigate the bias and variance of the empirical state-action visitation distribution $d_{\mathcal{D}}({\bm{s}},{\bm{a}})$ under PROPS, ROS, and on-policy sampling. We report the bias and variance averaged over all $({\bm{s}},{\bm{a}})\in{\mathcal{S}}\times{\mathcal{A}}$ computed as follows:
$\text{bias}=\frac{1}{|{\mathcal{S}}\times{\mathcal{A}}|}\sum_{({\bm{s}},{\bm{a}})\in{\mathcal{S}}\times{\mathcal{A}}}\left(\mathbb{E}\left[d_{{\mathcal{D}}}({\bm{s}},{\bm{a}})\right]-d_{\pi_{\bm{\theta}}}({\bm{s}},{\bm{a}})\right)$ (13)
$\text{variance}=\frac{1}{|{\mathcal{S}}\times{\mathcal{A}}|}\sum_{({\bm{s}},{\bm{a}})\in{\mathcal{S}}\times{\mathcal{A}}}\mathbb{E}\left[\left(d_{{\mathcal{D}}}({\bm{s}},{\bm{a}})-d_{\pi_{\bm{\theta}}}({\bm{s}},{\bm{a}})\right)^{2}\right]$ (14)
As shown in Fig. 8, the visitation distributions under PROPS and ROS empirically have near zero bias (note that the vertical axis has scale $10^{-18}$) and have lower variance than on-policy sampling.
Figure 10: Sampling error ablations with a fixed, random target policy. Here, “no clipping” refers to setting $\epsilon_{\text{PROPS}}{}=\infty$, and “no regularization” refers to setting $\lambda=0$. Solid curves denote the mean over 10 seeds, and shaded regions denote 95% bootstrap confidence intervals.
@@ D.2 Correcting Sampling Error for a Fixed Target Policy
In this appendix, we expand upon results presented in Section 6.2 of the main paper and provide additional experiments investigating the degree to which PROPS reduces sampling error with respect to a fixed, randomly initialized target policy. We additionally include ablation studies investigating the effects of clipping and regularization.
We tune PROPS and ROS using a hyperparameter sweep. For PROPS, we sweep over learning rates in $\{10^{-3},10^{-4}\}$ and fix the remaining PROPS hyperparameters: regularization coefficient $\lambda=0.1$, target KL $\delta_{\text{PROPS}}=0.03$, and clipping coefficient $\epsilon_{\text{PROPS}}=0.3$. For ROS, we sweep over learning rates in $\{10^{-3},10^{-4},10^{-5}\}$. We report results for hyperparameters yielding the lowest sampling error.
In Fig. 9, we see that PROPS achieves lower sampling error than both ROS and on-policy sampling across all tasks. ROS shows little to no improvement over on-policy sampling, again highlighting the difficulty of applying ROS to higher dimensional tasks with continuous actions.
Fig. 10 ablates the effects of PROPS’s clipping mechanism and regularization on sampling error reduction. We ablate clipping by setting $\epsilon_{\text{PROPS}}{}=\infty$, and we ablate regularization by setting $\lambda=0$. We use a fixed expert target policy and use the same tuning procedure described earlier in this appendix. In all tasks, PROPS achieves higher sampling error without clipping or regularization than it does with clipping and regularization, though this method nevertheless outperforms on-policy sampling in all tasks. Only removing clipping increases sampling error in most setups, and only removing regularization often increases sampling error for smaller batches of data e.g., 1024 samples. These observations indicate that while regularization is helpful, clipping has a stronger effect on sampling error reduction than regularization when the target policy is fixed.
@@ D.3 Correcting Sampling Error During RL Training
In this appendix, we include additional experiments investigating the degree to which PROPS reduces sampling error during RL training, expanding upon results presented in Section 6.3 of the main paper. We include sampling error curves for all six MuJoCo benchmark tasks and additionally provide ablation studies investigating the effects of clipping and regularization on sampling error reduction and data efficiency in the RL setting. We ablate clipping by tuning RL agents with $\epsilon_{\text{PROPS}}{}=\infty$, and we ablate regularization by tuning RL agents with $\lambda=0$. Fig. 11 and Fig. 12 show sampling error curves without clipping and without regularization, respectively. Without clipping, PROPS achieves larger sampling error than on-policy sampling in all tasks except Humanoid. Without regularization, PROPS achieves larger sampling error in 3 out of 6 tasks. These observations indicate that while clipping and regularization both help reduce sampling error during RL training, clipping has a stronger effect on sampling error reduction. As shown in Fig. 13 PROPS data efficiency generally decreases when we remove clipping or regularization.
Lastly, we consider training with larger buffer sizes $b$ in Fig. 14. We find that data efficiency may decrease with a larger buffer size. Intuitively, the more historic data kept around, the more data that must be collected to impact the aggregate data distribution.
Figure 13: Mean return over 50 seeds of PROPS with and without clipping or regularizing the PROPS objective. Shaded regions denote 95% bootstrap confidence intervals.
Figure 14: Mean return over 50 seeds for PROPS with different buffer sizes. We exclude $b=8$ for Humanoid-v4 due to the expense of training and tuning. Shaded regions denote 95% bootstrap confidence intervals.
Appendix E Runtime Comparisons
@@
Figure 15 shows runtimes for PROPS, PPO-Buffer, and PPO averaged over 3 runs. We trained all agents on a MacBook Air with an M1 CPU and used the same tuned hyperparameters used throughout the paper. PROPS takes at most twice as long as PPO-Buffer; intuitively, both PROPS and PPO-Buffer learn from the same amount of data but PROPS learns two policies.
We note that PPO-Buffer is faster than PPO in HalfCheetah-v4 because, with our tuned hyperparameters, PPO-Buffer performs fewer target policy updates than PPO. In particular, PPO-Buffer is updating its target policy every 4096 steps, whereas PPO is updating the target policy every 1024 steps.
Appendix F Hyperparameter Tuning for RL Training
For all RL experiments in Section 6.3 and Appendix D.3, we tune PROPS, PPO-Buffer, and PPO separately using a hyperparameter sweep over parameters listed in Table 2 and fix the hyperparameters in Table 4 across all experiments. Since we consider a wide range of hyperparameter values, we ran 10 independent training runs for each hyperparameter setting. We then performed 50 independent training runs for the hyperparameter settings yielding the largest returns at the end of RL training. We report results for these hyperparameters in the main paper.
PPO learning rate $10^{-3},10^{-4}$, linearly annealed to 0 over training
@@ Appendix F Hyperparameter Tuning for RL Training
PROPS behavior update frequency $m$ $256,512,1024$
PROPS KL cutoff $\delta_{\text{PROPS}}$ $0.03,0.05,0.1$
PROPS regularizer coefficient $\lambda$ $0.1,0.3$
@@
PPO PPO PROPS PROPS PROPS PROPS
Environment Batch Size Learning Rate Update Freq. Learning Rate KL Cutoff Regularization $\lambda$
Swimmer-v4 2048 $10^{-3}$ 1024 $10^{-5}$ 0.03 0.1
Hopper-v4 2048 $10^{-3}$ 256 $10^{-3}$ 0.05 0.3
HalfCheetah-v4 1024 $10^{-4}$ 512 $10^{-3}$ 0.05 0.3
Walker2d-v4 2048 $10^{-3}$ 256 $10^{-3}$ 0.1 0.3
Ant-v4 2048 $10^{-4}$ 256 $10^{-3}$ 0.03 0.1
Humanoid-v4 8192 $10^{-4}$ 256 $^{-4} 0.1 0.1$
Table 3: Hyperparameters selected from our hyperparameter sweep for RL training.
@@
Buffer size $b$ 2 target batches (also 3, 4, and 8 in Fig. 14)
@@
PPO and PROPS optimizers Adam (Kingma and Ba, 2015)
@@
Table 4: Hyperparameters fixed across all experiments. We use the PPO implementation provided by CleanRL (Huang et al., 2022).
Appendix G Hyperparameter Sensitivity Analysis
To assess how sensitive PROPS is to hyperparameter selection, we follow Patterson et al. (2023) and plot RL training performance aggregated over a large subset of runs from our hyperparameter sweep. Although we include the PPO target learning rate and PPO target batch size in our hyperparameter sweep, we keep them fixed at their tuned values in our violin plots. Including these hyperparameters would tell us about the sensitivity of PPO, but we are only interested in understanding the sensitivity of PROPS. We run 10 seeds for each hyperparameter setting. Fig. 17 and Fig. 17 show training curves and performance profiles for tuned PROPS and aggregated PROPS as well as our baselines. Across tasks, tuned PROPS achieves higher returns than aggregated PROPS, demonstrating the benefit of careful hyperparameter tuning. Nevertheless, aggregated PROPS still outperforms PPO and PPO-Buffer, suggesting that PROPS maintains a degree of hyperparameter robustness.
Figure 16: Mean return over 50 seeds. Shaded regions denote 95% confidence intervals.
Figure 17: Performance profiles over 50 seeds. Shaded regions denote 95% confidence intervals.